OpenStudy (anonymous):

what is the x intercept of y=1+(1/x)+(1/x^2)

OpenStudy (anonymous):

Is the x intercept when y = 0?

OpenStudy (anonymous):

Yess!

OpenStudy (anonymous):

Then shouldn't you just try to solve \[0=1 + \frac{1}{x}+\frac{1}{x^2}\]?

OpenStudy (anonymous):

I did & i dont think theres a defined answer so i was just checking if i did something wrong

OpenStudy (anonymous):

Ah, did you check the discriminant to see if it was positive/zero/negative?

OpenStudy (cwrw238):

multiply each term by x^2

OpenStudy (cwrw238):

giving x^2 + x + 1 = 0 and check discriminant as freewill said

OpenStudy (cwrw238):

yup

OpenStudy (anonymous):

alright thanks.

OpenStudy (cwrw238):

the roots are non-real (complex) numbers

OpenStudy (cwrw238):

yw